OpenStudy (anonymous):

Suppose f(1)=3, f'(1)=2, g(2)=2, g'(2)=6, h(x)=g(2x) and f(x)=f(x)/g(2x) find h'(1) and f'(1)

OpenStudy (anonymous):

\[ f'(x)=\frac{f'(x)g(2x)-2f(x)g(2x)}{(g(2x))^2}\] \[\huge f'(x)=2g(2x) \] h'(1)=2*2=4 can u do the first one f'(x)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I meant to write h'(x)=2g(2x) the first one is given by you f'(1)=2

OpenStudy (anonymous):

those answers aren't correct, not working on my homework